# Diffusion Cocktail: Fused Generation from Diffusion Models

### v1

@@ Abstract
Diffusion models excel at generating high-quality images and are easy to extend, making them extremely popular among active users who have created an extensive collection of diffusion models with various styles by fine-tuning base models such as Stable Diffusion. Recent work has focused on uncovering semantic and visual information encoded in various components of a diffusion model, enabling better generation quality and more fine-grained control. However, those methods target improving a single model and overlook the vastly available collection of fine-tuned diffusion models. In this work, we study the combinations of diffusion models. We propose Diffusion Cocktail (Ditail), a training-free method that can accurately transfer content information between two diffusion models. This allows us to perform diverse generations using a set of diffusion models, resulting in novel images that are unlikely to be obtained by a single model alone. We also explore utilizing Ditail for style transfer, with the target style set by a diffusion model instead of an image. Ditail offers a more detailed manipulation of the diffusion generation, thereby enabling the vast community to integrate various styles and contents seamlessly and generate any content of any style.
Figure 1: Diffusion Cocktail demo use cases. Our method primarily focuses on transferring content information between diffusion models. For the top-left example, our method can manipulate the generation process of multiple diffusion models by injecting source image structure, resulting in novel generated images of various styles. We can extend the idea to the other three use cases, including style transfer on real images, stylized regional content injection, and stylized image editing with prompts.
@@ 1 Introduction
Figure 2: Comparisons of various diffusion-based image manipulation settings. Unlike all previous methods, Ditail influences the generation of images by incorporating information encoded in multiple DMs. With Ditail, the user can benefit from the largely available set of DMs of dramatically different styles/contents and accurate transfer information between them to get the desirable generation.
Diffusion models (DMs) have shown great success in generating high-quality images [35, 12, 3], conditioning on user inputs such as text snippets [28, 24, 27, 16, 11], reference images [22] or enriched text [6]. A critical feature of the current DMs is the ease of extending and fine-tuning (e.g., using ControlNet [40] or LoRA [13]). The strong capability in image quality and the easy-to-extend feature of DMs result in a vast community of non-industrial/non-academic users fine-tuning the base DMs to generate images of specific styles or specializing certain characters of interest  [^1] To date there are about 7K user-generated DMs checkpoints and about 68K LoRAs available on Civitai: https://civitai.com/.. This gives us access to an extensive collection of fine-tuned DMs with similar model architectures but diverse contents and styles.
@@
Even though much work has focused on improving a single DM, rarely have people studied combinations of diffusion models, ignoring the vast collection of DMs created by everyday users. To the best of our knowledge, we are the first to study image generation using multiple DMs and the transfer of information between diffusion models. We propose Diffusion Cocktail (Ditail) that is readily applicable to existing DMs without further training/fine-tuning, thus allowing us to utilize the abundant and fast-growing DM resources efficiently and effectively.
Ditail enables us to accurately transfer the content/structure information between two DMs during the diffusion process. Given two DMs with different styles (e.g., one fine-tuned with watercolor images and the other fine-tuned with line art), in addition to the two images generated by the two DMs individually, Ditail creates two new images, each of which exhibits the style of one DM while complying with the structure of the other DM. This allows users to first generate images using $m$ DMs independently, identify images with desirable contents and styles, and apply Ditail to mix the content and styles for more satisfying results. Ditail can, in principle, generate $\mathcal{O}(m^{2})$ novel images that are not obtainable using a single model from the existing $m$ DMs, which greatly increases the generation flexibility and diversity. DMs with specialized styles or content could get more applications by incorporating other models using Ditail.
Moreover, Ditail is extensible to the style transfer of an input image that is not necessarily generated by a DM. We propose a novel style transfer task: the target style is set according to a generative model, namely a style model, as opposed to a target style image in the classic style transfer setting. Given a reference image and a target style DM, we first extract the content information of the reference image using a general-purpose base DM model (e.g., Stable Diffusion 1.5) and then apply Ditail to transfer style from the base model to the target style model. If no target style model is available, we can fine-tune a DM using a set of style images, thereby essentially composing a mixed target style based on this set of images. At the cost of collecting more images and fine-tuning the model, we benefit from richer information about the target style that no single image could provide.
Finally, Ditail allows more fine-grained control over the process and uncovers the various DM component properties. Ditail could create a smooth interpolation between the content information of the two DMs. We consider both positive and negative prompts in Ditail and provide parameters to control the content corresponding to each prompt, thus creating more diverse generation results. Ditail could also transfer partial content information by applying sub-masks on the self-attention and feature maps in U-Net, empowering users to mix the spatial structures of reference images.
@@ 2 Related Work
(a) Overview of the Ditail framework
(b) U-Net Layer
Figure 3: Overview of Ditail for style transfer. The content image first goes through diffusion inversion, ideally using a DM matching the content image’s style. The resulting latents then get injected into the style DM for generation with optional controls such as masking and guidance strength. Given various-style DMs, we generate images with their styles while preserving the content image’s structures. Note if the content image is DM-generated, we may skip the inversion process as we have access to the latent information (novel generation).
@@
Image manipulation with diffusion models. Pix2pix-zero [25] propose an image-to-image translation method that preserves the source image content without manual prompting to guide the translation. It automatically discovers editing directions that reflect desired edits in the text embedding space and retains the source image content by applying a cross-attention guidance loss throughout the diffusion process.  Hertz et al. [8] identify the importance of cross-attention in controlling the spatial information and its mapping to the text prompt, thus leading to three types of image editing. A few other works have attempted to manipulate image generation with personalized object images as guidance [30, 31, 5]. Orthogonally, our work aims to enable personalized manipulation of image styles through the choice of target style DM, a realistic setup with many fine-tuned DMs available. Tumanyan et al. [37] propose a plug-and-play (PnP) method that edits a given image using a DM with extra prompts to guide the manipulation. PnP first performs diffusion inversion of the given image and extracts features and self-attention maps to record the structure and semantic information. Inspired by PnP, Ditail generalizes the concept of PnP to transfer information between two DMs and facilitates downstream tasks.
Inversion of diffusion models. The inversion problem of a DM aims to find the noise maps as well as the conditioning vector that could generate a given image during the reverse diffusion process. Null-text inversion [23] enables intuitive text-based editing of real images with Stable Diffusion models by using an initial DDIM inversion as an anchor, which only tunes the null-text embedding used in classifier-free guidance. On the other hand, the inversion of the deterministic DDIM sampling process [3] and the inversion of the CLIP [27] have been studied to derive a latent noise map for generating specific real images. Gal et al. [5] propose a textual inversion technique for discovering a new pseudo-word describing the visual concept of a user-provided concept with optimization on only a few images.
@@ 3 Method
Preliminaries. Diffusion models [3, 12, 28, 33] are probabilistic generative models, whose generation consists of the forward and the backward process. The forward process is defined as a Markov chain where each step adds a small amount of Gaussian noise to the latent variables. Formally, for a sequence of time steps $t=1,\dots,T$, the process is:
@@
Stable Diffusion [28], a pre-trained text-conditioned Latent Diffusion Model (LDM), has gained attention for its ability to generate high-quality images from textual descriptions. It adopts a pre-trained image autoencoder [17] to perform the diffusion processes in the latent space with an extra condition on the text embedding ${\boldsymbol{c}}$. It uses a U-Net [29] architecture, where each layer consists of a residual block, a self-attention block, and a cross-attention block (Figure 2(b)). At some time step $t$, we denote the output of the $l$-th U-Net block as ${\boldsymbol{z}}_{t}^{l}$; the residual block output as ${\boldsymbol{f}}_{t}^{l}$; the projected self-attention features as ${\boldsymbol{q}}_{t}^{l},{\boldsymbol{k}}_{t}^{l},{\boldsymbol{v}}_{t}^{l}$.
Figure 4: Comparison of content injection. The middle column refers to content injection by replacing ${\boldsymbol{f}}_{t},{\boldsymbol{q}}_{t}$, and ${\boldsymbol{k}}_{t}$. The right column refers to our proposed injection via decoding ${\boldsymbol{z}}_{t}$.
@@ 3.1 Ditail Content Injection
Inspired by previous work studying the guidance and translation of DM [14, 38, 37], we transfer the content information from one DM to another by fusing the target DM U-Net hidden states with the ones from the source DM.
@@
We have two optional control parameters for Ditail. The classifier-free guidance $\omega$ [11] increases the fidelity of the generated image to the positive prompt. The mask $\mathcal{M}$ selects the subset of the latent representations for transfer so that only a portion of the content gets injected. Unlike previous work that injects the features (${\boldsymbol{f}}_{t}$) and self-attention keys and queries (${\boldsymbol{q}}_{t},{\boldsymbol{k}}_{t}$), we inject the content information via decoding the source latent ${\boldsymbol{z}}_{t}$. Comparing the two, injection of ${\boldsymbol{z}}_{t}$ greatly reduces the space required to store the injection information (i.e., from 5.7Gb to 4.7Mb in the case of 50 editing steps). Empirically, we find no significant difference between the two approaches qualitatively (Fig. 4).
@@
6:$\tilde{{\boldsymbol{c}}}\leftarrow\text{concat}({\boldsymbol{n}},{\boldsymbol{e}}_{pos},{\boldsymbol{e}}_{neg})$
@@
9:    ${\boldsymbol{z}}_{t}\leftarrow\text{concat}({\boldsymbol{z}}^{src}_{t},{\boldsymbol{z}}^{tgt}_{t},{\boldsymbol{z}}^{tgt}_{t})$
10:    ${\boldsymbol{\hat{\epsilon}}}_{t-1}^{pos},{\boldsymbol{\hat{\epsilon}}}_{t-1}^{neg}\leftarrow\text{noise-pred-with-inj}\!\left({\boldsymbol{z}}_{t},\tilde{{\boldsymbol{c}}},\mathcal{M},t;\theta^{tgt}\right)$
11:    ${\boldsymbol{\hat{\epsilon}}}_{t-1}\leftarrow{\boldsymbol{\hat{\epsilon}}}_{t-1}^{neg}+\omega({\boldsymbol{\hat{\epsilon}}}_{t-1}^{pos}-{\boldsymbol{\hat{\epsilon}}}_{t-1}^{neg})$
12:    ${\boldsymbol{z}}^{tgt}_{t-1}\leftarrow\text{DDIM-samp}({\boldsymbol{z}}^{tgt}_{t},{\boldsymbol{\hat{\epsilon}}}_{t-1})$
13:end for
@@
3.2 Ditail for Novel Generation
The Ditail content injection allows us to generate new images by mixing the representations of two DMs. As described in Alg. 2, given a pair of DMs, we input the positive and negative prompts to generate two images from each DM individually. During the generation process, we can also store the latent representation $\{{\boldsymbol{z}}_{t}^{(1)}\},\{{\boldsymbol{z}}_{t}^{(2)}\}$, and we inject the latent of the first DM to the generation of the second DM and vice versa. By doing so, we generate two more images, each consisting of the structure information of one DM and the style information of the other DM.
Given a set of $m$ DMs, we can use Ditail to generate $m^{2}$ images as opposed to the $m$ images by running every DM independently. This dramatically expands the generation capacity, producing images that are not attainable through a single DM. It also allows the user to have better control over the generation process as well as the integration of more domain-specific DMs. The user may first use some DMs to generate specific content structure and then transfer the spatial information to other DMs with desirable styles. Fig. 5 shows the cross-transfer among three DMs.
3.3 Ditail for Model Style Transfer
As Ditail excels at injecting content information into the diffusion process while preserving the individual DM’s style, we can also treat Ditail as a method for style transfer between models. Essentially, we copy the spatial features and self-attention maps from the content DM and inject those into the style DM. A more precise description for this procedure could be “content transfer” instead of style transfer, as the style DM, which received content information from the other DM, is responsible for generating the target image.
The typical style transfer task takes two images as inputs, and the target is to transfer the style from a style image to the content image while preserving the spatial structure of the content image. Suppose the content image is generated using a DM of the same architecture as the style DM, and we have access to the injection features/self-attention maps; then, we can directly apply style transfer using Ditail as illustrated in the diverse generation. On the contrary, given an arbitrary content image (real or generated), we can first extract the structure information using diffusion inversion and then inject such information using Ditail (Alg. 3).
Algorithm 2 Ditail for Novel Generation
1:Input: Prompts $(P_{pos},P_{neg})$, two DMs $\theta^{(1)},\theta^{(2)}$
2:Output: Two generated images and two novel images
3:$\mathcal{I}^{(1)},\{{\boldsymbol{z}}_{t}^{(1)}\}_{t=1:T}\leftarrow\mbox{DM}(P_{pos},P_{neg};\theta^{(1)})$
4:$\mathcal{I}^{(2)},\{{\boldsymbol{z}}_{t}^{(2)}\}_{t=1:T}\leftarrow\mbox{DM}(P_{pos},P_{neg};\theta^{(2)})$
5:$\mathcal{I}^{(1)\rightarrow(2)}\leftarrow\mbox{Ditail}(\{{\boldsymbol{z}}_{t}^{(1)}\}_{t=1:T},P_{pos},P_{neg};\theta^{(2)})$
6:$\mathcal{I}^{(2)\rightarrow(1)}\leftarrow\mbox{Ditail}(\{{\boldsymbol{z}}_{t}^{(2)}\}_{t=1:T},P_{pos},P_{neg};\theta^{(1)})$
7:return $\mathcal{I}^{(1)},\mathcal{I}^{(2)},\mathcal{I}^{(1)\rightarrow(2)},\mathcal{I}^{(2)\rightarrow(1)}$
Figure 5: Demonstration of style transfer across different DMs. The diagonal images are the non-modified images generated using each diffusion model with the same textual prompt. Every column corresponds to the images generated by one diffusion model, using contents from other models.
Algorithm 3 Ditail for Style Transfer

### v2

# Diffusion Cocktail: Mixing Domain-Specific Diffusion Models for Diversified Image Generations

@@ Abstract
Diffusion models, capable of high-quality image generation, receive unparalleled popularity for their ease of extension. Active users have created a massive collection of domain-specific diffusion models by fine-tuning base models on self-collected datasets. Recent work has focused on improving a single diffusion model by uncovering semantic and visual information encoded in various architecture components. However, those methods overlook the vastly available set of fine-tuned diffusion models and, therefore, miss the opportunity to utilize their combined capacity for novel generation. In this work, we propose Diffusion Cocktail (Ditail), a training-free method that transfers style and content information between multiple diffusion models. This allows us to perform diversified generations using a set of diffusion models, resulting in novel images unobtainable by a single model. Ditail also offers fine-grained control of the generation process, which enables flexible manipulations of styles and contents. With these properties, Ditail excels in numerous applications, including style transfer guided by diffusion models, novel-style image generation, and image manipulation via prompts or collage inputs.
Figure 1: Diffusion Cocktail demo use cases. Ditail primarily focuses on transferring content information between diffusion models. We control the generation process of multiple diffusion models by injecting source image structure, resulting in novel images of various contents and styles. This core idea can be extended to various use cases. Better view with color. Zoom in for best view.
@@ 1 Introduction
Diffusion models (DMs) have shown great success in generating high-quality images [37, 12, 4], conditioning on user inputs such as text snippets [30, 26, 29, 16, 13], reference images [24] or enriched text [7]. A critical feature of the current DMs is the ease of extending and fine-tuning (e.g., ControlNet [42] or LoRA [14]). The strong capability in image quality and the easy-to-extend feature of DMs result in a vast community of non-industrial/non-academic users fine-tuning the base DMs to generate images of specific styles or specializing certain characters of interest  [^1] To date, there are about 8K user-generated DM checkpoints and about 100K LoRAs available on Civitai: https://civitai.com/.. This gives us access to a vast collection of fine-tuned DMs with similar model architectures but diverse contents and styles.
@@
Even though much work has focused on improving a single DM, rarely have people studied the combinations of diffusion models, ignoring the vast collection of DMs created by everyday users. To the best of our knowledge, we are the first to study image generation using multiple DMs and the transfer of information in between. We propose Diffusion Cocktail (Ditail), a simple yet effective method that is readily applicable to existing DMs without further training/fine-tuning, thus allowing us to utilize the abundant and fast-growing DM resources efficiently and effectively.
Ditail enables us to transform the source image (that is not necessarily generated by a DM) to some target domains, using domain-specific DM checkpoints as proxies. By accurately transferring the content information between two DMs during the diffusion process, Ditail can, in theory, handle any-to-any transformation as long as we have DMs fine-tuned on certain domains. In other words, we explore a novel model-centric style transfer paradigm, where the target style is set according to a generative model, as opposed to a target style image in the conventional style transfer setting. Given a source image and a target style DM (with optional LoRA), we first extract the content information from DDIM inversion [36] and then apply Ditail to transform the image to the target domain via the target style DM. If no target style model is available, one can mix existing DM and LoRA checkpoints to tailor a novel target style or fine-tune a new checkpoint using a set of style images (essentially defining a target domain based on this set of images). In this way, we can benefit from robust stylization and manipulation capability that no single image could provide.
Moreover, Ditail is highly extensible to various side applications, such as novel image generation and content manipulation. Given $m$ images generated by $m$ DMs independently with the same textual prompt, one can identify images with desirable contents and styles, and apply Ditail to mix the content and styles for more satisfying results. Ditail can, in principle, generate $\mathcal{O}(m^{2})$ novel images that are not obtainable using a single model from the existing $m$ DMs, significantly increasing the generation flexibility and diversity. Besides, Ditail can also be used to manipulate the contents by either modifying the textual prompts (e.g., changing the color of the watch in Fig. 2(a)) or aligning the appearance of a collage (e.g., adding a pair of sunglasses to an image in Fig. 2(b)).
Finally, Ditail allows more fine-grained control over the process and uncovers the DM component properties. We consider both positive and negative prompts in Ditail for content manipulation and further propose a novel condition scaling technique to control the extent of content preservation, thus creating more diversified generation results.
@@ 2 Related Work
Image manipulation with a diffusion model. Pix2pix-zero [27] propose an image-to-image translation method that preserves the source image content without manual prompting to guide the translation. It automatically discovers editing directions that reflect desired edits in the text embedding space and retains the source image content by applying a cross-attention guidance loss throughout the diffusion process. Hertz et al. [9] identify the importance of cross-attention in controlling the spatial information and its mapping to the text prompt, thus leading to three types of image editing. A few other works have attempted to manipulate image generation with personalized object images as guidance [32, 33, 6]. Orthogonally, our work aims to enable personalized manipulation of image styles through the choice of target style DM, a realistic setup with many fine-tuned DMs available. Tumanyan et al. [39] propose a plug-and-play (PnP) method that edits a given image using a DM with extra prompts to guide the manipulation. PnP first performs diffusion inversion of the given image and extracts features and self-attention maps to record the structure and semantic information. Inspired by PnP, Ditail generalizes the concept of PnP to transfer information between two DMs and facilitates downstream tasks.
Inversion of a diffusion model. The inversion problem of a DM aims to find the noise maps as well as the conditioning vector that could generate a given image during the reverse diffusion process. Null-text inversion [25] enables intuitive text-based editing of real images with Stable Diffusion models by using an initial DDIM inversion as an anchor, which only tunes the null-text embedding used in classifier-free guidance. On the other hand, the inversion of the deterministic DDIM sampling process [4] and the inversion of the CLIP [29] have been studied to derive a latent noise map for generating specific real images. Gal et al. [6] propose a textual inversion technique for discovering a new pseudo-word describing the visual concept of a user-provided concept with optimization on only a few images.
Figure 2: (a) Ditail for model-centric style transfer and content manipulation. The content image first goes through diffusion inversion, ideally using a DM matching the content image’s domain. The resulting latents are then fed into the style DM for denoising and content injection. Given various-style DMs, we generate images with their styles while preserving the content image’s structures; (b) Ditail can be naturally extended to transform collages to the same target style/domain; (c) Ditail allows tailoring novel target styles by merging existing style DMs via an interpolation hyper-parameter $\gamma$. Better view with color. Zoom in for best view.
@@ 3 Method
3.1 Preliminaries
Diffusion models [4, 12, 30, 35] are probabilistic generative models, whose generation consists of the forward and the backward process. The forward process is defined as a Markov chain where each step adds a small amount of Gaussian noise to the latent variables. Formally, for a sequence of time steps $t=1,\dots,T$, the process is:
@@ 3.1 Preliminaries
Stable Diffusion [30], a pre-trained text-conditioned Latent Diffusion Model (LDM), has gained attention for its ability to generate high-quality images from textual descriptions. It adopts a pre-trained image autoencoder [17] to perform the diffusion processes in the latent space with an extra condition on the text embedding ${\boldsymbol{c}}$. It uses a U-Net [31] architecture, where each layer consists of a residual block, a self-attention block, and a cross-attention block. At time step $t$, we denote the output of the U-Net block as ${\boldsymbol{z}}_{t}$; the residual block outputs as $\{{\boldsymbol{h}}_{t}^{l}\}$; the projected self-attention features as $\{{\boldsymbol{q}}_{t}^{l}\},\{{\boldsymbol{k}}_{t}^{l}\},\{{\boldsymbol{v}}_{t}^{l}\}$, where $l$ denotes layer index.
@@ 3.2 Ditail Content Injection
Inspired by previous work studying the guidance and translation of DM [15, 40, 39], we transfer the content information from one DM to another by fusing the U-Net hidden states of the target DM with the ones from the source DM.
@@
We offer two optional control parameters in Ditail. The classifier-free guidance $\omega$ [13] increases the fidelity of the generated image to the positive prompt. The mask $\mathcal{M}$ selects the subset of the latent representations for transfer so that only a portion of the content gets injected. Unlike previous work that injects self-attention maps and features from cached tensors, we inject the content information along with the noise sampling of ${\boldsymbol{z}}_{t}$, where $\{{\boldsymbol{q}}_{t}^{l}\},\{{\boldsymbol{k}}_{t}^{l}\},\{{\boldsymbol{h}}_{t}^{l}\}$ can be reconstructed from ${\boldsymbol{z}}_{t}^{src}$ due to the deterministic nature of DDIM [36]. In other words, Line 7-8 of Alg. 1 can be merged to perform content injection and noise prediction in one operation (see more details in Appendix). Comparing the two, the proposed method greatly reduces the space required to store the injection information (i.e., from 5.7 GB to 4.7 MB in the case of 50 inference steps). One potential caveat here is that we are using different DM checkpoints for inversion and sampling, which may fail to reconstruct the original hidden states; however, our quantitative results over 300 test images and 5 DMs indicate that the performance drops are low enough to be negligible, possibly due to the fact that most DM checkpoints, despite the various styles they exhibit, still share similar weight distributions in the U-Net. Hence, we adopt the proposed content injection paradigm by default to be more memory-friendly and efficient (i.e., it only requires one noise sampling process instead of two).
@@
6:$\tilde{{\boldsymbol{c}}}\leftarrow\text{concat}({\boldsymbol{e}}_{pos},{\boldsymbol{e}}_{neg})$
@@
9:    ${\boldsymbol{z}}_{t}\leftarrow\text{concat}({\boldsymbol{z}}^{tgt}_{t},{\boldsymbol{z}}^{tgt}_{t})$
10:    $\{{\boldsymbol{q}}_{t}^{l}\},\{{\boldsymbol{k}}_{t}^{l}\},\{{\boldsymbol{h}}_{t}^{l}\}\leftarrow\text{noise-pred}({\boldsymbol{z}}^{src}_{t},{\boldsymbol{n}},t;\theta^{tgt})$
@@
12:                        $\left({\boldsymbol{z}}_{t},\tilde{{\boldsymbol{c}}},\mathcal{M},t;\{{\boldsymbol{q}}_{t}^{l}\},\{{\boldsymbol{k}}_{t}^{l}\},\{{\boldsymbol{h}}_{t}^{l}\};\theta^{tgt}\right)$
13:    ${\boldsymbol{\hat{\epsilon}}}_{t-1}\leftarrow{\boldsymbol{\hat{\epsilon}}}_{t-1}^{neg}+\omega({\boldsymbol{\hat{\epsilon}}}_{t-1}^{pos}-{\boldsymbol{\hat{\epsilon}}}_{t-1}^{neg})$
14:    ${\boldsymbol{z}}^{tgt}_{t-1}\leftarrow\text{DDIM-samp}({\boldsymbol{z}}^{tgt}_{t},{\boldsymbol{\hat{\epsilon}}}_{t-1})$
15:end for
[…]
3.3 Ditail for Model-Centric Style Transfer
As Ditail excels at injecting content information into the diffusion process while preserving the individual DM’s domain specialization, we can also treat Ditail as a method for style transfer between models. Essentially, we copy the spatial features and self-attention maps from the content DM and inject those into the style DM. A more precise description for this procedure could be “content transfer” instead of style transfer, as the style DM, which received content information from the other DM, generates the target image.
The typical style transfer task takes two images as inputs, and the target is to transfer the style from a style image to the content image while preserving the spatial structure of the content image. Suppose the content image is generated using a DM of the same architecture as the style DM, and we have access to the features/self-attention maps; then, we can directly apply style transfer using Ditail. On the contrary, as illustrated in Fig. 2(a), given an arbitrary content image (real or generated), we can first extract the structure information using diffusion inversion (i.e., adding reversely predicted noise to the latents using DDIM [36]) and then inject such information using Ditail (Alg. 2).
Algorithm 2 Ditail for Model-Centric Style Transfer and Content Manipulation